# Learning Grouped Lattice Vector Quantizers for Low-Bit LLM Compression

### v1

@@ Abstract
footnotetext: # Corresponding author.
@@ 1 Introduction
Large language models (LLMs) have achieved extraordinary success across a wide range of natural language processing tasks, including machine translation, question answering, text generation, and summarization vaswani2017attention ; devlin2019bert ; radford2019language . These models, exemplified by the GPT radford2019language , BERT devlin2019bert , Llama touvron2023llama , Qwen bai2023qwen and DeepSeek liu2024deepseek , are at the forefront of AI innovation. Beyond supervised pretraining, reinforcement learning techniques, most notably reinforcement learning from human feedback (RLHF), have further enhanced LLMs by aligning them with human preferences and improving their safety, helpfulness, and controllability stiennon2020learning ; ouyang2022training ; bai2022training ; liu2025survey ; liu2025verigui ; fang2025serl . However, their exceptional performance comes at exorbitant computation and memory costs, which hinders their deployment on edge devices and in other resource-limited scenarios wang2021survey . For instance, state-of-the-art models with billions of parameters often require hundreds of gigabytes of memory and expensive hardware for inference brown2020language ; radford2019language , posing serious challenges for scalability and accessibility.
Quantization has emerged as a key technique to alleviate the resource burdens by reducing the precision of model weights from 16-bit or 32-bit floating-point representations to low-bit integers hubara2017quantizing ; banner2018post . This significantly reduces the memory footprint and computation load, improving model economy and inference speed. A relatively simple method is Post-training quantization (PTQ), which quantizes a pre-trained LLM model without retraining  li2021brecq ; frantar2022gptq . However, PTQ has difficulties in achieving ultra-low bit-widths, such as 2 or 3 bits per weight dettmers2022gpt3 . Severe quantization errors can lead to significant degradation in model performance, such as higher perplexity in language tasks frantar2022gptq or noticeable drops in downstream quality, thereby limiting the practical utility of low-bit quantization.
@@
Recently, some progresses have been made in LLM weight quantization, including vector quantization (VQ) techniques tseng2024quip ; egiazarian2024extreme ; tseng2024qtip and salience-aware bit allocation huang2024slim . Vector quantization methods, such as QuIP# tseng2024quip , leverage structured codebooks (e.g., E8 lattice) to improve quantization fidelity by aligning the codebook structure with the statistical properties of model weights. However, QuIP# is constrained by the use of fixed lattice designs across the entire model, which fails to account for the diversity of parameter distributions across different groups or layers. This uniform approach can lead to suboptimal quantization for certain groups, ultimately limiting both compression performance and overall model quality. In contrast, AQLM egiazarian2024extreme proposes learning free-form VQs for different groups, which allows for more flexible quantization. However, this approach has the drawback that decoding requires the lookup operation, which is computationally more expensive than QUIP# and other existing PTQ methods.
@@
In this paper, we propose a novel framework, Grouped Lattice Vector Quantization (GLVQ) (illustrated in Fig. 1), aiming to increase the LLM compression ratio by reducing quantization resolution to be below 3 bits per weight. Specifically, GLVQ end-to-end optimizes the geometric structure of the lattice codebook for each quantization group based on that group’s unique parameter distribution. By shaping the lattice to fit the data distribution of each group, we reduce quantization distortion, preserve critical information better, and improve model fidelity over nonadaptive lattices under extreme compression regimes. Compared to free-form VQ used in AQLM egiazarian2024extreme , GLVQ employs a structured codebook, hence it enjoys the operational advantage of much lower quantization complexity. Further investigations reveal that LLM parameter distributions can be highly skewed or heavy-tailed. If this unfortunately happens, optimizing GLVQ alone has limited effect, because the lattice cells of each group have congruent shape, albeit optimized, and have regular configurations. To counteract this, we incorporate group-specific companding transformations. These functions reshape each group’s parameter distribution into a more uniform space before applying the lattice quantizer, thereby reducing distortion in the low-magnitude regions and enabling the lattice to allocate codepoints more effectively. Together, group-specific lattice codebook optimization and companding form a powerful synergy that maintains high accuracy while significantly reducing bit-width requirements.
@@
We validate our approach on popular LLM benchmarks, demonstrating its effectiveness in improving perplexity and inference efficiency compared to state-of-the-art quantization methods frantar2022gptq ; tseng2024quip ; egiazarian2024extreme ; ding2023cbq . Our results show that GLVQ not only enables higher compression ratios but also maintains robust performance across diverse language tasks, making it a practical solution for deploying LLMs in resource-constrained environments. Furthermore, our framework provides insights into the interplay between parameter distribution and quantization fidelity, opening new directions for research in efficient model compression.
@@ 2.1 LLM Post-Training Quantization (PTQ)
Recent surveys gholami2022survey ; gong2024survey have emphasized the importance of post-training quantization (PTQ) methods in the context of large language models (LLMs), particularly as model scales reach billions of parameters. Classic approaches such as AdaRound nagel2020up , BitSplit wang2020towards , AdaQuant hubara2021accurate , BRECQ li2021brecq , and OBQ frantar2022optimal have been shown to work well for smaller networks. However, applying these techniques directly to LLMs often incurs prohibitive computational overhead due to their reliance on expensive solvers and per-layer fine-grained adjustments.
@@
In an effort to address scalability challenges, multiple PTQ frameworks have emerged. For instance, GPT3.int8() dettmers2022gpt3 , ZeroQuant yao2022zeroquant , and Lut-gemm park2022lut employed straightforward round-to-nearest (RTN) quantization strategies and introduced granular control to manage the trade-off between model size and accuracy. GPTQ frantar2022gptq proposed a more data-aware layer-wise optimization to minimize the $\ell_{2}$ reconstruction error. Follow-up studies dettmers2023case argued that 4-bit RTN might be near-optimal for certain classes of networks, but data-aware schemes like GPTQ can exceed this limit without compromising efficiency. Meanwhile, 8-bit quantization of both weights and activations has also been investigated dettmers2022gpt3 ; xiao2023smoothquant ; yao2022zeroquant , revealing that “outlier features” can substantially degrade performance, and spurring specialized outlier-handling techniques.
@@
Subsequent efforts prioritized addressing such outliers in weight matrices. For example, SpQR dettmers2023spqr maintains a sparse, higher-precision representation of large-magnitude weights to mitigate their outsized influence. Similarly, AWQ lin2024awq employs per-channel scaling to handle channels with large activation magnitudes, while SqueezeLLM kim2023squeezellm adopts diagonal Fisher approximations and K-means clustering to perform non-uniform quantization, selectively preserving critical weights in full precision. A notable result is QuIP chee2024quip , which strategically rotates (e.g., via Hadamard transform) and then projects weights onto a lattice to control outliers, enabling 2-bit per-parameter compression with only minor increases in perplexity.
@@
Beyond these developments, recent research introduces vector quantization (VQ) to further enhance LLM compression. Transformer-VQ lingle2023transformer applies VQ to the Attention key vectors, effectively reducing Attention complexity to linear time. Other VQ-based approaches generally optimize discrete code assignments (often in the form of one-hot vectors) and their associated codebooks egiazarian2024extreme ; tseng2024quip . For instance, AQLM egiazarian2024extreme proposes an additive quantization strategy that is input-adaptive, while QuIP# tseng2024quip leverages a highly symmetric $E_{8}$ lattice to handle weights that follow (approximately) a spherical sub-Gaussian distribution. GPTVQ van2024gptvq interleaves column-wise quantization steps with updates to the unquantized parameters, guided by the Hessian of the per-layer output reconstruction MSE, and then employs SVD and integer quantization for additional codebook compression. PV-Tuning malinovskii2024pv shows how relying on straight-through estimators (STE) alone can be suboptimal, proposing a stage-wise method that separately optimizes continuous parameters (e.g., scales, codebooks) and discrete assignments. QTIP tseng2024qtip decouples the notion of codebook size from both bitrate and dimensionality through a stateful decoding, thereby facilitating ultra-high-dimensional VQ. NestQuant savkin2025nestquant proposes a novel PTQ scheme for weights and activations that is based on self-similar nested lattices.
@@ 2.2 Lattice Vector Quantization (LVQ)
Scalar quantization maps each parameter to a discrete code individually. Vector quantization, instead, quantizes a group of parameters as a vector by mapping it to the nearest code vector from a codebook Gray1984 . Lattice-based quantization conway1987sphere ; gibson1988lattice ; ErezZamir2004 goes a step further, leveraging a structured grid in the high-dimensional space for efficient quantization. Formally, let $\mathbf{G}\in\mathbb{R}^{d\times d}$ be a generation matrix, which defines a full-rank lattice: $\Lambda=\bigl\{\mathbf{G}\,\mathbf{z}\,\big|\,\mathbf{z}\in\mathbb{Z}^{d}\bigr\}.$ Each column of $\mathbf{G}$ acts as a basis vector in $\mathbb{R}^{d}$, so any lattice point can be written as an integer combination of these basis vectors.
@@
In general, solving the nearest–lattice–point problem is both computationally prohibitive for large $d$ and non-differentiable, motivating the use of efficient approximate methods such as Babai rounding Babai1986 . Once the integer lattice index $\hat{\mathbf{z}}$ is found and stored, we can decode the quantized vector by multiplying back: $\hat{\mathbf{x}}\;=\;\mathbf{G}\,\hat{\mathbf{z}}$, yielding a point on the lattice that approximates the original vector $\mathbf{x}$.
Learnable LVQ. Recent studies have shown that learning or adapting the generation matrix $\mathbf{G}$ to a given data distribution can significantly reduce quantization error at a fixed bit budget khalil2023ll ; zhang2024learning ; xu2025multirate ; xu2025improving ; zhang2025receptive . This is especially beneficial in low-bit regimes where per-dimension scalar methods often struggle. Additionally, practical implementations of LVQ can incorporate companding or pre/post-transformation steps to further align the data with the lattice geometry ErezZamir2004 ; zhang2023lvqac .
@@ 2.3 Companding for Quantization
Many neural network weight distributions exhibit high dynamic range and non-uniform behavior (Gaussian or sub-Gaussian), which can amplify quantization errors in the low-magnitude regime. Companding (derived from “compressing” and “expanding”) smith1957instantaneous ; kaneko1970unified ; tsividis1990companding offers a simple, yet effective strategy to reshape these distributions before quantization, thereby reducing distortion in critical regions. Typically, the companding process comprises two stages:
@@
A well-known example of $F$ is the $\mu$-law transformation ITU-T_G711 ; Haykin1994 , originally used in audio signal processing:
@@ 3 Method
    Lattice codebook learning. With $b_{g}$ fixed (implying a lattice codebook of $2^{b_{g}}$ points), we reshape every weight group $\mathbf{W}_{g}\in\mathbb{R}^{m_{g}\times n_{g}}$ into a matrix of size $d\times\ell_{g}$ ($d$ is the lattice dimension). We then learn a group-specific generation matrix $\mathbf{G}_{g}\in\mathbb{R}^{d\times d}$ and the corresponding integer indices $\mathbf{Z}_{g}\in\mathbb{Z}^{d\times\ell_{g}}$ can be obtained via Babai rounding Babai1986 . Hence the quantized weight matrix is
@@ 3.1 Salience-Determined Bit Allocation
To achieve optimal bit allocation under strict resource constraints, we adopt the Salience-Determined Bit Allocation (SDBA) mechanism proposed in Slim-LLM huang2024slim , which dynamically assigns bit-widths to each group based on its importance.
@@ 3.2 Lattice Codebook Learning
The key innovation behind GLVQ is the ability to adaptively learn a specialized lattice structure for each weight group, rather than imposing a fixed, universal lattice (such as the $E_{8}$ lattice employed by QuIP# tseng2024quip ). Recent studies, such as Slim-LLM huang2024slim , show that even after standard preprocessing, large language models (LLMs) exhibit significant heterogeneity across weight groups, motivating the need for group-specific quantization schemes. Specifically, GLVQ learns a generation matrix $\mathbf{G}_{g}\in\mathbb{R}^{d\times d}$ for each weight group $g$, where the dimension $d$ (e.g., chosen from $\{8,16,32,\dots\}$) controls both computational complexity and the granularity of quantization. Crucially, the integer lattice indices $\mathbf{Z}_{g}$ are not directly optimized, but rather determined via Babai rounding, making $\mathbf{G}_{g}$ the only explicitly learned parameter.
@@
    Lattice index assignment (fix $\mathbf{G}_{g}$). Fixing the lattice basis $\mathbf{G}_{g}$, we update integer indices using Babai rounding Babai1986 :

### v2

@@ Abstract
footnotetext: Corresponding author.
@@ 1 Introduction
Large language models (LLMs) have achieved extraordinary success across a wide range of natural language processing tasks, including machine translation, question answering, text generation, and summarization Vaswani et al. (2017); Devlin et al. (2019); Radford et al. (2019). These models, exemplified by the GPT Radford et al. (2019), BERT Devlin et al. (2019), Llama Touvron et al. (2023), Qwen Bai et al. (2023) and DeepSeek Liu et al. (2024), are at the forefront of AI innovation. Beyond supervised pretraining, reinforcement learning techniques, most notably reinforcement learning from human feedback (RLHF), have further enhanced LLMs by aligning them with human preferences and improving their safety, helpfulness, and controllability Stiennon et al. (2020); Ouyang et al. (2022); Bai et al. (2022); Liu et al. (2025a, b); Fang et al. (2025). However, their exceptional performance comes at exorbitant computation and memory costs, which hinders their deployment on edge devices and in other resource-limited scenarios Cheng et al. (2017). For instance, state-of-the-art models with billions of parameters often require hundreds of gigabytes of memory and expensive hardware for inference Brown et al. (2020); Radford et al. (2019), posing serious challenges for scalability and accessibility.
Quantization has emerged as a key technique to alleviate the resource burdens by reducing the precision of model weights from 16-bit or 32-bit floating-point representations to low-bit integers Krishnamoorthi (2018); Liu et al. (2021). This significantly reduces the memory footprint and computation load, improving model economy and inference speed. A relatively simple method is Post-training quantization (PTQ), which quantizes a pre-trained LLM model without retraining  Li et al. (2021); Frantar et al. (2022). However, PTQ has difficulties in achieving ultra-low bit-widths, such as 2 or 3 bits per weight Dettmers et al. (2022). Severe quantization errors can lead to significant degradation in model performance, such as higher perplexity in language tasks Frantar et al. (2022) or noticeable drops in downstream quality, thereby limiting the practical utility of low-bit quantization.
@@
Recently, some progresses have been made in LLM weight quantization, including vector quantization (VQ) techniques Tseng et al. (2024a); Egiazarian et al. (2024); Tseng et al. (2024b) and salience-aware bit allocation Huang et al. (2024). Vector quantization methods, such as QuIP# Tseng et al. (2024a), leverage structured codebooks (e.g., E8 lattice) to improve quantization fidelity by aligning the codebook structure with the statistical properties of model weights. However, QuIP# is constrained by the use of fixed lattice designs across the entire model, which fails to account for the diversity of parameter distributions across different groups or layers. This uniform approach can lead to suboptimal quantization for certain groups, ultimately limiting both compression performance and overall model quality. In contrast, AQLM Egiazarian et al. (2024) proposes learning free-form VQs for different groups, which allows for more flexible quantization. However, this approach has the drawback that decoding requires the lookup operation, which is computationally more expensive than QUIP# and other existing PTQ methods.
@@
In this paper, we propose a novel framework, Grouped Lattice Vector Quantization (GLVQ) (illustrated in Fig. 1), aiming to increase the LLM compression ratio by reducing quantization resolution to be below 3 bits per weight. Specifically, GLVQ end-to-end optimizes the geometric structure of the lattice codebook for each quantization group based on that group’s unique parameter distribution. By shaping the lattice to fit the data distribution of each group, we reduce quantization distortion, preserve critical information better, and improve model fidelity over nonadaptive lattices under extreme compression regimes. Compared to free-form VQ used in AQLM Egiazarian et al. (2024), GLVQ employs a structured codebook, hence it enjoys the operational advantage of much lower quantization complexity. Further investigations reveal that LLM parameter distributions can be highly skewed or heavy-tailed. If this unfortunately happens, optimizing GLVQ alone has limited effect, because the lattice cells of each group have congruent shape, albeit optimized, and have regular configurations. To counteract this, we incorporate group-specific companding transformations. These functions reshape each group’s parameter distribution into a more uniform space before applying the lattice quantizer, thereby reducing distortion in the low-magnitude regions and enabling the lattice to allocate codepoints more effectively. Together, group-specific lattice codebook optimization and companding form a powerful synergy that maintains high accuracy while significantly reducing bit-width requirements.
@@
We validate our approach on popular LLM benchmarks, demonstrating its effectiveness in improving perplexity and inference efficiency compared to state-of-the-art quantization methods Frantar et al. (2022); Tseng et al. (2024a); Egiazarian et al. (2024); Ding et al. (2023). Our results show that GLVQ not only enables higher compression ratios but also maintains robust performance across diverse language tasks, making it a practical solution for deploying LLMs in resource-constrained environments. Furthermore, our framework provides insights into the interplay between parameter distribution and quantization fidelity, opening new directions for research in efficient model compression.
@@ 2.1 LLM Post-Training Quantization (PTQ)
Recent surveys Gholami et al. (2022); Gong et al. (2025) have emphasized the importance of post-training quantization (PTQ) methods in the context of large language models (LLMs), particularly as model scales reach billions of parameters. Classic approaches such as AdaRound Nagel et al. (2020), BitSplit Wang et al. (2020), AdaQuant Hubara et al. (2021), BRECQ Li et al. (2021), and OBQ Frantar and Alistarh (2022) have been shown to work well for smaller networks. However, applying these techniques directly to LLMs often incurs prohibitive computational overhead due to their reliance on expensive solvers and per-layer fine-grained adjustments.
@@
In an effort to address scalability challenges, multiple PTQ frameworks have emerged. For instance, GPT3.int8() Dettmers et al. (2022), ZeroQuant Yao et al. (2022), and Lut-gemm Park et al. (2022) employed straightforward round-to-nearest (RTN) quantization strategies and introduced granular control to manage the trade-off between model size and accuracy. GPTQ Frantar et al. (2022) proposed a more data-aware layer-wise optimization to minimize the $\ell_{2}$ reconstruction error. Follow-up studies Dettmers and Zettlemoyer (2023) argued that 4-bit RTN might be near-optimal for certain classes of networks, but data-aware schemes like GPTQ can exceed this limit without compromising efficiency. Meanwhile, 8-bit quantization of both weights and activations has also been investigated Dettmers et al. (2022); Xiao et al. (2023); Yao et al. (2022), revealing that “outlier features” can substantially degrade performance, and spurring specialized outlier-handling techniques.
@@
Subsequent efforts prioritized addressing such outliers in weight matrices. For example, SpQR Dettmers et al. (2023) maintains a sparse, higher-precision representation of large-magnitude weights to mitigate their outsized influence. Similarly, AWQ Lin et al. (2024) employs per-channel scaling to handle channels with large activation magnitudes, while SqueezeLLM Kim et al. (2023) adopts diagonal Fisher approximations and K-means clustering to perform non-uniform quantization, selectively preserving critical weights in full precision. A notable result is QuIP Chee et al. (2023), which strategically rotates (e.g., via Hadamard transform) and then projects weights onto a lattice to control outliers, enabling 2-bit per-parameter compression with only minor increases in perplexity.
@@
Beyond these developments, recent research introduces vector quantization (VQ) to further enhance LLM compression. Transformer-VQ Lingle (2023) applies VQ to the Attention key vectors, effectively reducing Attention complexity to linear time. Other VQ-based approaches generally optimize discrete code assignments (often in the form of one-hot vectors) and their associated codebooks Egiazarian et al. (2024); Tseng et al. (2024a). For instance, AQLM Egiazarian et al. (2024) proposes an additive quantization strategy that is input-adaptive, while QuIP# Tseng et al. (2024a) leverages a highly symmetric $E_{8}$ lattice to handle weights that follow (approximately) a spherical sub-Gaussian distribution. GPTVQ Van Baalen et al. (2024) interleaves column-wise quantization steps with updates to the unquantized parameters, guided by the Hessian of the per-layer output reconstruction MSE, and then employs SVD and integer quantization for additional codebook compression. PV-Tuning Malinovskii et al. (2024) shows how relying on straight-through estimators (STE) alone can be suboptimal, proposing a stage-wise method that separately optimizes continuous parameters (e.g., scales, codebooks) and discrete assignments. QTIP Tseng et al. (2024b) decouples the notion of codebook size from both bitrate and dimensionality through a stateful decoding, thereby facilitating ultra-high-dimensional VQ. NestQuant Savkin et al. (2025) proposes a novel PTQ scheme for weights and activations that is based on self-similar nested lattices.
@@ 2.2 Lattice Vector Quantization (LVQ)
Scalar quantization maps each parameter to a discrete code individually. Vector quantization, instead, quantizes a group of parameters as a vector by mapping it to the nearest code vector from a codebook Gray (1984). Lattice-based quantization Conway and Sloane (2013); Gibson and Sayood (1988); Erez and Zamir (2004) goes a step further, leveraging a structured grid in the high-dimensional space for efficient quantization. Formally, let $\mathbf{G}\in\mathbb{R}^{d\times d}$ be a generation matrix, which defines a full-rank lattice: $\Lambda=\bigl\{\mathbf{G}\,\mathbf{z}\,\big|\,\mathbf{z}\in\mathbb{Z}^{d}\bigr\}.$ Each column of $\mathbf{G}$ acts as a basis vector in $\mathbb{R}^{d}$, so any lattice point can be written as an integer combination of these basis vectors.
@@
In general, solving the nearest–lattice–point problem is both computationally prohibitive for large $d$ and non-differentiable, motivating the use of efficient approximate methods such as Babai rounding Babai (1986). Once the integer lattice index $\hat{\mathbf{z}}$ is found and stored, we can decode the quantized vector by multiplying back: $\hat{\mathbf{x}}\;=\;\mathbf{G}\,\hat{\mathbf{z}}$, yielding a point on the lattice that approximates the original vector $\mathbf{x}$.
Learnable LVQ. Recent studies have shown that learning or adapting the generation matrix $\mathbf{G}$ to a given data distribution can significantly reduce quantization error at a fixed bit budget Khalil et al. (2023); Zhang and Wu (2024); Xu et al. (2025b, a); Zhang and Wu (2025). This is especially beneficial in low-bit regimes where per-dimension scalar methods often struggle. Additionally, practical implementations of LVQ can incorporate companding or pre/post-transformation steps to further align the data with the lattice geometry Erez and Zamir (2004); Zhang and Wu (2023).
@@ 2.3 Companding for Quantization
Many neural network weight distributions exhibit high dynamic range and non-uniform behavior (Gaussian or sub-Gaussian), which can amplify quantization errors in the low-magnitude regime. Companding (derived from “compressing” and “expanding”) Smith (1957); Kaneko (1970); Tsividis et al. (1990) offers a simple, yet effective strategy to reshape these distributions before quantization, thereby reducing distortion in critical regions. Typically, the companding process comprises two stages:
@@
A well-known example of $F$ is the $\mu$-law transformation Recommendation (1988); Bhagyaveni et al. (2022), originally used in audio signal processing:
@@ 3 Method
    Lattice codebook learning. With $b_{g}$ fixed (implying a lattice codebook of $2^{b_{g}}$ points), we reshape every weight group $\mathbf{W}_{g}\in\mathbb{R}^{m_{g}\times n_{g}}$ into a matrix of size $d\times\ell_{g}$ ($d$ is the lattice dimension). We then learn a group-specific generation matrix $\mathbf{G}_{g}\in\mathbb{R}^{d\times d}$ and the corresponding integer indices $\mathbf{Z}_{g}\in\mathbb{Z}^{d\times\ell_{g}}$ can be obtained via Babai rounding Babai (1986). Hence the quantized weight matrix is
@@ 3.1 Salience-Determined Bit Allocation
To achieve optimal bit allocation under strict resource constraints, we adopt the Salience-Determined Bit Allocation (SDBA) mechanism proposed in Slim-LLM Huang et al. (2024), which dynamically assigns bit-widths to each group based on its importance.
@@ 3.2 Lattice Codebook Learning
The key innovation behind GLVQ is the ability to adaptively learn a specialized lattice structure for each weight group, rather than imposing a fixed, universal lattice (such as the $E_{8}$ lattice employed by QuIP# Tseng et al. (2024a)). Recent studies, such as Slim-LLM Huang et al. (2024), show that even after standard preprocessing, large language models (LLMs) exhibit significant heterogeneity across weight groups, motivating the need for group-specific quantization schemes. Specifically, GLVQ learns a generation matrix $\mathbf{G}_{g}\in\mathbb{R}^{d\times d}$ for each weight group $g$, where the dimension $d$ (e.g., chosen from $\{8,16,32,\dots\}$) controls both computational complexity and the granularity of quantization. Crucially, the integer lattice indices $\mathbf{Z}_{g}$ are not directly optimized, but rather determined via Babai rounding, making $\mathbf{G}_{g}$ the only explicitly learned parameter.
@@
    Lattice index assignment (fix $\mathbf{G}_{g}$). Fixing the lattice basis $\mathbf{G}_{g}$, we update integer indices using Babai rounding Babai (1986):